Two forces act on an object. The first force has a magnitude of 17.0 N and is oriented 58.0� counter-clockwise from the x axis, while the 2nd force has x- and y-components (-21.0 N, 10.5 N). Express the magnitude and direction of the net force.
Answer #1

x-component of first force=17 cos 58=9N

y-component of first force=17 Sin 58=14.42 N

Net force=-21i+9i+10.5j+14.42j=-12i+24.92j

Magnitude=27.66 N

Direction=180- Tan-1(24.92/12)=180-64.29=115.71 degrees counterclockwise from the horizontal